MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2007/05/10
OBSERVER: Sumi Miyake Paul Matsubara
CONDITION: fine -> slightly cloudy -> cloudy, sometimes windy
COMMENT: D2249-60: OK. D2250-300: OK. Dark noise for a vacume sensor (hot cathod + pyrani) was tested at D2251-D2253. The current broken vacuume sensor was replaced by a repaired one.
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 18:38 END 06:51
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 6360-6387 | 28 | 154 | 9240 |
LMC | 1517-1521 | 5 | 25 | 7500 |
SMC | 371-371 | 1 | 2 | 600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 34 | 181 | 17340 (04h49m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
7 | 7 | 7 | 7 | 28 | 5 | 6 | 6 | 26 | 5 | 5 | 5 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
4 | 4 | 4 | 4 | 4 | 4 | 4 | 4 | 4 | 4 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 1 | 1 | 5 | 5 | 4 | 1 | 1 | 1 |
Field13 | Field14 | Field15 | Field16 |
1 | 1 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CHECK SHEET
Time | Status | Weather |
18:00 | observing | fine |
19:00 | observing | fine |
20:00 | observing | fine |
21:00 | observing | fine |
22:00 | observing | slightly cloudy |
23:00 | observing | slightly cloudy |
00:00 | observing | cloudy |
01:00 | waiting | windy |
02:00 | observing | fine, slightly windy |
03:00 | waiting | cloudy |
04:00 | waiting | cloudy |
05:00 | waiting | cloudy |
06:00 | waiting | cloudy |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day